The width of rectangle is 25% shorter than the length. If the perimeter is 21. Find the are of the rectangle


1
Expert's answer
2021-11-30T11:35:09-0500

"P=2l+2w=21."

"w=(1-0.25)l."

"2l+2(1-0.25)l=21."

"2l+1.5l=21."

"l=\frac{21}{3.5}=6."

"w=0.75l=6*0.75=4.5."

"A=lw=6*4.5=27."